OpenStudy (anonymous):

3x + 5y=3 x+2y=0 solve

OpenStudy (nikato):

Elimination or substitution?

OpenStudy (anonymous):

elimination

OpenStudy (anonymous):

wait no the question says substitution

OpenStudy (nikato):

Ok, subtract 2y on both sides in the second equation

OpenStudy (anonymous):

x= -2y + 0

OpenStudy (nikato):

So x=-2y Substitute this value of x into the first equation, so... 3(-2y)+5y=3 And u should be able to solve for y

OpenStudy (anonymous):

-y = 3?

OpenStudy (nikato):

Yes. And multiply both sides by -1 so y=-3. Plug this value of y into one of the original equation to solve for x

OpenStudy (anonymous):

the second one or first

OpenStudy (nikato):

Doesn't matter. Whichever one i find easier

OpenStudy (nikato):

You*

OpenStudy (anonymous):

-6y + x = 0 ?

OpenStudy (anonymous):

so x = -6y

OpenStudy (nikato):

Are u plugging it into the second equation ?

OpenStudy (anonymous):

yea

OpenStudy (nikato):

Should be x+2(-3)=0 And now solve for x

OpenStudy (anonymous):

x=6

OpenStudy (anonymous):

so (6,-3)

OpenStudy (nikato):

Correct